OpenStudy (anonymous):

Working together, 7 identical pumps can empty a pool in 6 hours. How many hours will it take 4 pumps to empty the same pool?

OpenStudy (anonymous):

@satellite73 can we take this one slow too?

OpenStudy (anonymous):

sure unit rate again

OpenStudy (anonymous):

So 7/6 will be how much 1 pump can empty a pool

OpenStudy (anonymous):

careful

OpenStudy (anonymous):

lets start with how much 7 pumps do per hour

OpenStudy (anonymous):

isn't that what 7/6 finds?

OpenStudy (anonymous):

7 pumps take 6 hours, so 7 pumps do \(\frac{1}{6}\) per hour

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

how much does one pump do? don't multiply by 7, divide (like last time)

OpenStudy (anonymous):

1/42

OpenStudy (anonymous):

ok good now each pump does \(\frac{1}{42}\) per hour, how much does 4 pumps do (this time multiply)

OpenStudy (anonymous):

4/42

OpenStudy (anonymous):

right

OpenStudy (anonymous):

or you can reduce to get \(\frac{2}{21}\)

OpenStudy (anonymous):

hmmmm did we do something wrong? my answer say 42/4 = 10 1/2

OpenStudy (anonymous):

you want to see how long it takes to do one job,

OpenStudy (anonymous):

solve \[\frac{2}{21}T=1\] making \[T=\frac{21}{2}\]

OpenStudy (anonymous):

no we didn't do anything wrong, we just were not done yet we got the unit rate at \(\frac{2}{21}\) but we need the total time

OpenStudy (anonymous):

ohhh okay

OpenStudy (anonymous):

they do 2/21 of a job per hour, takes 21/2 hours to do one job

OpenStudy (anonymous):

ohhh okay

OpenStudy (anonymous):

there may have been a quicker way to do this, but it is good to know how to get the unit rate so you can do all kinds of problems with it

OpenStudy (anonymous):

hmmm ok

OpenStudy (anonymous):

so do we now multiply by 4?

OpenStudy (anonymous):

no now we are done

OpenStudy (anonymous):

ohhh okay this is a different method i got it now

OpenStudy (anonymous):

we got the unit rate of \(\frac{1}{42}\) that was pump per hour

OpenStudy (anonymous):

then we multiplied by 4 to get the rate for 4 pumps and got \(\frac{2}{21}\)

OpenStudy (anonymous):

ahhhh ic

OpenStudy (anonymous):

then we solved \[\frac{2}{21}T=1\] the 1 being one pool

OpenStudy (anonymous):

if it was two pools it would have been \[\frac{2}{21}=2\]

OpenStudy (anonymous):

hmmm okay i think i understand now i'll review the last 2 problems we did a couple of times

OpenStudy (anonymous):

ok if you get stuck let me know

OpenStudy (anonymous):

i will thanks <333

OpenStudy (anonymous):

we can make one up and do it now if you like

OpenStudy (anonymous):

yes yes that would help even more

OpenStudy (anonymous):

see that they are not that hard and are largely the same

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

you want me to make it up, or you want to?

OpenStudy (anonymous):

can u ?

OpenStudy (anonymous):

k they numbers might not come out real nice if i make it up off the top of my head, but no matter

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

it takes 5 hours for 8 men to build one wall how long would it take 10 men to build 2 walls (sexist i now but whatever)

OpenStudy (anonymous):

first find the unit rate for one man let me know when you think you have it

OpenStudy (anonymous):

1/8 divided by 5 = 1/40

OpenStudy (anonymous):

yay!

OpenStudy (anonymous):

now the rate for 10 men

OpenStudy (anonymous):

1/40 * 10 = 10/40 = 1/4

OpenStudy (anonymous):

ok good now how long to build 2 walls

OpenStudy (anonymous):

1/4 * 2= 2/4= 1/2

OpenStudy (anonymous):

aw damn we were doing so well

OpenStudy (anonymous):

oh crap...

OpenStudy (anonymous):

not going to take half an hour !

OpenStudy (anonymous):

now we need to go slow either we solve \[\frac{1}{4}T=2\]

OpenStudy (anonymous):

ohhhhhh yeaaaa thats right so 8

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

ok i'll keep practicing this

OpenStudy (anonymous):

almost got it 100% thanks satellite

OpenStudy (anonymous):

ok, i think it is getting easier though right? first part went quick

OpenStudy (anonymous):

yw